Delia Landscaping is considering a new 4-year project. The necessary fixed assets will cost $195,000 and be depreciated on a 3-year MACRS and have no salvage value. The MACRS percentages each year are 33.33 percent, 44.45 percent, 14.81 percent, and 7.41 percent, respectively. The project will have annual sales of $132,000, variable costs of $35,500, and fixed costs of $12,850. The project will also require net working capital of $3,450 that will be returned at the end of the project. The company has a tax rate of 35 percent and the project's required return is 12 percent. What is the net present value of this project?
Answer #1

NPV FIXED ASSESTS WORKING CAPITAL year OCF -195,000.00 -3,450.00 disou cnt FCF rate discounted 12% value -198,450.00 1.00 -19

257 258 calculation of net profit year TAX OCF 259 260 261 net profit before dep 1 83,650.00 2 83,650.00 3 83,650.00 4 83,650

237 NPV 238 WORKING CAPITAL year FIXED ASSESTS OCF 239 240 0 -195000 - 3450 disoucnt FCF rate 0.12 =+B240+D240+C240 1 =+B241+

B 258 calculation of net profit year TAX OCF 259 260 1 261 2 2623 263 4. 264 net profit before dep =132000-35500-12850 =13200In second year there is no tax value as theere is loss. the depreciation amount is more than the profit value.

please ask if any query